The score of driving test has a normal distribution with mean 70 if given the standard deviation of sample is eight. A driving school's instructor claimed that if the candidate learned more than three hours per week, the mean score would be different than 70. A driving test was given to a random sample of 50 candidates with the mean score was 78.

a) state the null and alternative hypothesis
b) identify the type I error and type II error that correspond to the hypothesis above
c) test the claim at 5% level of significance
1
Expert's answer
2020-11-23T14:49:28-0500

"a)\ H_0: a=a_0=70, H_1: a\neq a_0=70\\
\overline{x}=78\\
N=50\\
s=8\\
c)\ \alpha=0.05\\
\text{We will use the following random variable:}\\
T=\frac{(\overline{X}-a_0)\sqrt{n}}{S}\\
T\text{ has t-distribution with } k=n-1\text{ degrees of freedom}.\\
\text{Observed value:}\\
t_{obs}=\frac{(78-70)\sqrt{50}}{8}\approx 7.07\\
\text{Critical value:}\\
t_{cr}=t_{cr}(\alpha;k)\approx 2.0096\\
\text{Critical region: }(-\infty,-2.0096)\cup (2.0096,\infty)\\
t_{obs}\text{ falls into the critical region. So we reject } H_0.\\
b)\text{ Type I Error}=\alpha=0.05.\\
\text{We will find Type II Error }\beta.\\
\frac{(\overline{X}-70)\sqrt{50}}{8}> -2.0096\\
\overline{X}> 67.73\\
\frac{(\overline{X}-70)\sqrt{50}}{8}< 2.0096\\
\overline{X}< 72.27\\
\beta=P\{67.73<\overline{X}< 72.27|a=78\}=\\
=P\{\frac{(67.73-78)\sqrt{50}}{8}<T< \frac{(72.27-78)\sqrt{50}}{8}\}=\\
=P\{-9.08<T< -5.06\}\approx 0\text{ (we use t-distribution table here)}."